width is 10 feet less than length. the perimeter is 100 feet. how do i find the area?
Answer by stanbon(48558)

width is 10 feet less than length. the perimeter is 100 feet. how do i find the area?
Let the length be "x".
Then the width is "x-10"
Perimeter=2(l+w)=100 ft.
l+w=50
x+x-10=50
2x=60
x=30 (length of the rectangle)
x-10=20 (width of the rectangle)
Area = 20*30=600 sq ft.
Cheers,
Stan H.